# DBConformer: Dual-Branch Convolutional Transformer for EEG Decoding

### v1

@@ IV-B Compared Approaches
Ten EEG decoding models were reproduced and compared with the proposed DBConformer.
@@ IV-C1 Evaluation Scenarios
To evaluate the performance under different conditions, we performed three experimental settings: chronological order (CO), cross-validation (CV), and leave-one-subject-out (LOSO), covering both within-subject and cross-subject settings, as in [17].
@@ IV-C2 Evaluation Metrics
Accuracy is employed to evaluate the performance on five MI datasets. Accuracy is the overall proportion of correctly classified trials, regardless of class:
$Accuracy=\frac{TP+TN}{TP+TN+FP+FN}\times 100\%,$ (13)
where TP is the number of correctly classified positive trials, TN the number of correctly classified negative trials, FP the number of negative trials misclassified as positive, and FN the number of positive trials misclassified as negative.
To evaluate the classification performance on class-imbalanced seizure datasets, three additional metrics are included.
  • Area under the receiver operating characteristic curve (AUC) is a threshold-independent metric that reflects the model’s ability to distinguish between seizure and non-seizure trials across varying decision thresholds.
  • F1 Score is the harmonic mean of precision and recall, indicating the balance between false positives and false negatives:
    $F1=\frac{2TP}{2TP+FP+FN}\times 100\%.$ (14)
  • Balanced classification accuracy (BCA) is the average of sensitivity and specificity, which accounts for class imbalance:
    $BCA=\frac{sensitivity+specificity}{2}\times 100\%,$ (15)
    where $sensitivity=\frac{TP}{TP+FN}$ is the proportion of correctly detected seizure trials, and $specificity=\frac{TN}{TN+FP}$ is the proportion of correctly detected normal trials.
@@ IV-C3 Parameter Settings
All experiments were repeated five times with a random seed list $[1,2,3,4,5]$, and the average results were reported. For all MI datasets, models were trained for 100 epochs using the Adam optimizer with a learning rate $10^{-3}$ and a batch size of 32. For seizure datasets, models are trained for 200 epochs using the Adam optimizer with a learning rate $10^{-3}$. The batch size was set to 256 for CHSZ and 512 for NICU because of the larger data size. Due to the specific characteristics of datasets, scenarios, and tasks, we adopted distinct model configurations accordingly, as summarized in Table IV.
TABLE IV: Model configurations of DBConformer for each dataset.
Dataset Patch Embed. # T-Trans. # T-Trans. # S-Trans. # S-Trans.
Size Size Layers Heads Layers Heads
CO/CV
BNCI2014001 125 40 2 2 2 2
BNCI2014004 125 40 2 2 2 2
Zhou2016 125 40 2 2 2 2
Blankertz2007 125 40 2 2 2 2
BNCI2014002 128 40 4 2 4 2
LOSO
BNCI2014001 125 40 2 2 2 2
BNCI2014004 125 40 2 2 2 2
Zhou2016 125 40 6 2 6 2
Blankertz2007 125 40 6 2 6 2
BNCI2014002 128 40 6 2 6 2
CHSZ 100 60 2 2 2 2
NICU 128 60 2 2 2 2
@@ IV-D Results on MI Classification
Table V presents the average classification accuracies of DBConformer and nine baseline models across five MI datasets under three evaluation settings. Observed that:
@@
  • Among the nine baseline models, IFNet achieves the highest average accuracy across all three settings, attributed to its frequency-aware architectural design. EEG Conformer ranks second under the CO and CV settings but exhibits reduced performance in the LOSO scenario, indicating that its large parameter size may cause overfitting when generalizing to unseen subjects.
@@
  • Compared to the CO setting, results under CV are generally higher across all models, consistent with findings in [17], where temporally ordered training limits generalization due to potential signal drift or session variance. CV benefits from more diverse training trials within each subject. This highlights the importance of fair and appropriate partitioning when benchmarking EEG decoding models.
@@
  • In the LOSO setting, DBConformer maintains the best generalization across subjects. This highlights the robustness of DBConformer in cross-subject EEG decoding tasks.
  • DBConformer consistently outperforms all baseline models across all datasets and settings. Compared to classical CNN models, such as EEGNet and FBCNet, DBConformer exhibits significant accuracy gains. Compared with recent CNN-Transformer hybrids (e.g., EEG Conformer and CTNet), DBConformer achieves consistently better results with fewer performance fluctuations. This demonstrates the effectiveness of the dual-branch design in capturing both temporal and spatial characteristics.
@@
TABLE V: Average classification accuracies (%) of DBConformer and nine baseline models on five MI datasets under CO, CV, and LOSO scenarios. The best average performance of each dataset is marked in bold, and the second best by an underline.
@@
Scenario Dataset EEGNet SCNN DCNN FBCNet ADFCNN CTNet EEG Conformer SlimSeiz IFNet DBConformer
CO BNCI2014001 69.05±1.00 73.57±2.36 59.29±1.64 68.97±1.26 73.73±2.26 73.49±2.05 78.57±0.66 68.89±2.05 77.94±0.93 79.05±1.82
BNCI2014004 68.43±2.20 68.15±1.39 62.41±1.64 65.46±1.57 69.63±1.19 71.57±1.33 72.04±1.57 68.24±2.74 73.43±1.72 75.09±1.53
Zhou2016 80.13±3.35 75.03±6.15 78.03±2.37 63.33±2.29 71.42±1.95 76.81±5.05 73.87±4.51 72.01±3.98 81.70±2.08 82.49±2.40
Blankertz2007 78.79±2.78 76.71±2.09 70.00±4.12 75.93±1.31 76.07±1.41 79.00±3.38 82.29±2.62 65.64±1.18 84.00±0.57 87.21±1.32
BNCI2014002 66.07±2.76 79.07±1.96 64.07±2.70 69.50±0.95 73.00±1.95 71.00±1.80 76.21±1.46 78.71±0.53 78.29±1.68 79.14±0.17
Average 72.49 74.51 66.76 68.64 72.77 74.38 76.59 70.70 79.07 80.60
CV BNCI2014001 71.86±0.96 78.22±0.19 61.59±1.08 74.74±0.67 77.27±0.60 75.97±0.76 83.62±0.99 75.56±1.48 82.62±0.91 83.66±0.40
BNCI2014004 68.91±0.86 68.13±0.63 63.51±0.40 67.17±0.30 69.48±1.24 70.96±1.06 73.15±0.72 69.74±0.46 72.02±0.52 74.17±0.63
Zhou2016 85.30±2.12 82.90±0.90 79.68±1.13 80.01±0.50 84.95±1.51 86.38±1.45 85.67±1.52 85.60±1.36 87.78±1.01 91.54±0.20
Blankertz2007 80.21±0.77 81.11±0.72 69.83±1.51 84.14±0.67 81.10±1.75 83.64±1.04 87.43±0.33 75.37±1.14 88.23±0.50 90.33±0.53
BNCI2014002 68.90±1.37 81.24±0.50 67.17±0.31 74.53±0.68 74.63±0.75 75.50±0.78 79.30±0.33 77.93±0.85 80.21±0.52 81.36±0.34
Average 75.04 78.32 68.36 76.12 77.49 78.49 81.83 76.84 82.17 84.21
LOSO BNCI2014001 73.64±1.14 72.22±1.03 73.21±1.79 72.56±0.96 71.76±0.75 73.40±1.22 73.07±2.01 69.62±1.32 74.52±0.75 77.67±0.61
BNCI2014004 67.78±0.92 62.19±0.79 62.56±0.56 67.09±0.69 65.17±0.66 65.28±1.09 64.22±1.05 63.15±1.42 67.69±0.43 69.85±0.47
Zhou2016 83.22±1.83 82.10±0.67 83.84±1.18 82.07±0.91 82.09±1.31 83.88±1.03 82.43±1.48 84.06±1.58 86.21±0.99 85.37±0.94
Blankertz2007 71.10±0.83 70.64±0.58 72.19±0.89 76.23±1.41 70.59±1.69 69.50±1.79 74.41±1.13 73.21±0.82 73.43±1.06 76.33±0.71
BNCI2014002 72.86±0.38 70.57±1.35 74.34±0.81 71.31±0.82 72.67±0.44 74.14±0.79 72.84±1.36 73.01±0.60 73.90±0.65 77.17±0.78
Average 74.54 71.54 73.23 73.85 72.45 73.24 73.40 72.61 75.15 77.28
@@ IV-E Results on Seizure Detection
For seizure, we compared the proposed DBConformer with seven baseline models: EEGNet, SCNN, DCNN, ADFCNN, EEG Conformer, SlimSeiz, and EEGWaveNet. The first five models were originally designed for tasks such as MI and emotion recognition, rather than seizure detection. But they are general for all EEG decoding tasks; thus, we re-implemented and adapted them to seizure detection. SlimSeiz [30] is a CNN-Mamba hybrid model for seizure prediction that can be easily implemented in the seizure detection task. EEGWaveNet [26] is a multi-scale depthwise CNN specifically developed for seizure detection, serving as a competitive task-specific benchmark. Table VI summarizes the average AUCs, F1 scores, and BCAs of DBConformer and six baseline models on the CHSZ and NICU datasets. Observed that:
@@
IV-F Ablation Study
To further validate the effectiveness of DBConformer’s key components, we conducted ablation studies on the parallel spatio-temporal design, positional encoding, and the channel attention module. Experiments were performed on five MI datasets under three settings, shown in Fig. 4. Observed that:
  • Parallel Spatio-Temporal Modeling. We compared the full DBConformer (T-Conformer + S-Conformer) with its T-Conformer only variant. Adding the S-Conformer branch consistently improves performance across all settings. Results confirm the effectiveness of incorporating spatial features via the auxiliary S-Conformer, which complements the dominant temporal modeling of T-Conformer.
  • Positional Encoding. Disabling the learnable positional embeddings for both temporal and spatial tokens results in performance drops under all settings. This validates the importance of explicit temporal/channel ordering when modeling patch and channel tokens.
  • Channel Attention Module. Replacing the channel attention module with naïve mean pooling degrades performance, confirming that adaptively reweighting EEG channels benefits enhancing spatial features.

### v2

@@ IV-B Compared Approaches
Thirteen EEG decoding models were compared with the proposed DBConformer in the experiments:
@@
  • MSCFormer [34] integrates multi-branch, multi-scale CNN modules with Transformer encoders to jointly model local and global EEG representations.
  • TMSA-Net [35] also integrates multi-scale CNNs with the local and global attention modules to enhance the feature extraction.
  • MSVTNet [36] extracts local spatio-temporal features at different filtered scales and introduces an auxiliary branch loss to facilitate effective CNN-Transformer integration.
@@ IV-C1 Evaluation Scenarios
To evaluate the performance under different conditions, we performed four experimental settings: chronological order (CO), cross-validation (CV), leave-one-subject-out (LOSO), and cross-dataset (CD), covering within-subject, cross-subject, and cross-device settings, as in [17, 63].
@@
  • CD: Cross-dataset validation, where training and testing were performed on distinct EEG datasets, e.g., training on BNCI2014001 and testing on BNCI2014004. The CD results are shown in Table S1 of the supplementary material.
@@ IV-C2 Evaluation Metrics
Accuracy is employed to evaluate the performance on MI and SSVEP. To evaluate the classification performance on class-imbalanced seizure datasets, three additional metrics, area under the receiver operating characteristic curve (AUC), weighted F11 score and balanced classification accuracy (BCA), are included in seizure detection.
@@ IV-C3 Parameter Settings
All experiments were repeated five times with a random seed list $[1,2,3,4,5]$, and the average results were reported. For all MI datasets, models were trained for 100 epochs using the Adam optimizer with a learning rate $10^{-3}$ and a batch size of 32. For seizure datasets, models are trained for 200 epochs using the Adam optimizer with a learning rate $10^{-3}$. The batch size was set to 256 for CHSZ and 512 for NICU because of the larger data size. For all baseline models, the model hyperparameters (e.g., kernel sizes, number of convolutional and Transformer layers) were set according to the original papers to ensure a fair comparison and reproducibility. Due to the specific characteristics of datasets, scenarios, and tasks, we adopted distinct model configurations accordingly, as summarized in Table S6 of the supplementary Material.
@@ IV-D Results on MI Classification
Table IV presents the average classification accuracies of DBConformer and nine baseline models across six MI datasets under three evaluation settings. Observe that:
@@
  • Among the nine baseline models, IFNet achieved the highest average accuracy across all three settings, attributed to its frequency-aware architectural design. EEG Conformer ranked second under the CO and CV settings but exhibits reduced performance in the LOSO scenario, indicating that its large parameter size may cause overfitting when generalizing to unseen subjects.
@@
  • Compared to the CO setting, results under CV were generally higher across all models, consistent with findings in [17], where temporally ordered training limits generalization due to potential signal drift or session variance. CV scenario benefited from more diverse training trials within each subject. This highlights the importance of fair and appropriate partitioning when benchmarking EEG decoding models.
@@
  • In the LOSO setting, DBConformer maintained the best generalization across subjects. This highlights the robustness of DBConformer in cross-subject EEG decoding tasks.
  • DBConformer almost always outperformed all 12 baseline models across all datasets and settings, demonstrating the effectiveness of the dual-branch design in capturing both temporal and spatial characteristics.
@@
TABLE IV: Average classification accuracies (%) of DBConformer and twelve baseline models on six MI datasets under CO, CV, and LOSO scenarios. The best average performance of each dataset is marked in bold, and the second best by an underline.
@@
Task Dataset EEGNet SCNN DCNN FBCNet ADFCNN CTNet EEG Confor. MSCFor. TMSA-Net MSVTNet SlimSeiz IFNet DBConfor.
CO 2014001 69.05±1.0 73.57±2.4 59.29±1.6 68.97±1.3 73.73±2.3 73.49±2.1 78.57±0.7 75.00±1.2 76.67±2.1 74.60±3.0 68.89±2.1 77.94±0.9 80.16±1.8
2014004 68.43±2.2 68.15±1.4 62.41±1.6 65.46±1.6 69.63±1.2 71.57±1.3 72.04±1.6 72.20±1.8 70.28±2.0 71.30±1.5 68.24±2.7 73.43±1.7 75.09±1.5
Zhou2016 80.13±3.4 75.03±6.2 78.03±2.4 63.33±2.3 71.42±2.0 76.81±5.1 73.87±4.5 73.93±4.5 67.94±3.0 70.32±1.4 72.01±4.0 81.70±2.1 82.49±2.4
Blankertz 78.79±2.8 76.71±2.1 70.00±4.1 75.93±1.3 76.07±1.4 79.00±3.4 82.29±2.6 85.29±1.6 78.93±1.0 84.29±0.9 65.64±1.2 84.00±0.6 87.21±1.3
2014002 66.07±2.8 79.07±2.0 64.07±2.7 69.50±1.0 73.00±2.0 71.00±1.8 76.21±1.5 75.79±1.9 76.79±1.9 78.93±1.4 78.71±0.5 78.29±1.7 79.14±0.2
OpenBMI 65.54±0.7 71.78±0.3 66.84±1.1 61.31±0.6 71.82±0.8 68.58±0.7 71.84±0.8 70.64±0.6 71.10±0.6 70.06±0.7 61.15±0.4 68.19±0.6 72.12±0.1
Average 71.33 74.05 66.77 67.42 72.61 73.41 75.80 75.47 73.62 74.91 69.11 77.26 79.37
CV 2014001 71.86±1.0 78.22±0.2 61.59±1.1 74.74±0.7 77.27±0.6 75.97±0.8 83.62±1.0 81.83±0.7 82.32±0.5 81.28±0.7 75.56±1.5 82.62±0.9 83.66±0.4
2014004 68.91±0.9 68.13±0.6 63.51±0.4 67.17±0.3 69.48±1.2 70.96±1.1 73.15±0.7 72.30±0.4 72.11±0.9 73.59±0.6 69.74±0.5 72.02±0.5 74.17±0.6
Zhou2016 85.30±2.1 82.90±0.9 79.68±1.1 80.01±0.5 84.95±1.5 86.38±1.5 85.67±1.5 89.16±0.8 85.54±1.3 88.06±1.2 85.60±1.3 87.78±1.0 91.54±0.2
Blankertz 80.21±0.8 81.11±0.7 69.83±1.5 84.14±0.6 81.10±1.8 83.64±1.0 87.43±0.3 88.41±0.9 74.09±0.6 89.07±0.8 75.37±1.1 88.23±0.5 90.33±0.5
2014002 68.90±1.4 81.24±0.5 67.17±0.3 74.53±0.7 74.63±0.8 75.50±0.9 79.30±0.3 80.29±0.7 80.41±0.6 80.91±0.5 77.93±0.9 80.21±0.5 81.36±0.3
OpenBMI 65.99±0.4 71.99±0.3 67.68±0.4 63.82±0.2 72.21±0.5 69.50±0.4 72.34±0.3 71.34±0.2 72.07±0.2 71.20±0.2 61.23±0.5 71.75±0.2 72.89±0.2
Average 73.53 77.27 68.24 74.07 76.61 76.99 80.25 80.55 77.76 80.69 74.24 80.44 82.32
LOSO 2014001 73.64±1.1 72.22±1.0 73.21±1.8 72.56±1.0 71.76±0.8 73.40±1.2 73.07±2.0 76.02±0.8 77.32±0.5 77.10±1.3 69.62±1.3 74.52±0.7 77.67±0.6
2014004 67.78±0.9 62.19±0.8 62.56±0.6 67.09±0.7 65.17±0.7 65.28±1.1 64.22±1.1 64.72±0.6 66.02±0.8 67.41±0.8 63.15±1.4 67.69±0.4 69.85±0.5
Zhou2016 83.22±1.8 82.10±0.7 83.84±1.2 82.07±0.9 82.09±1.3 83.88±1.0 82.43±1.5 85.31±2.4 80.78±2.3 84.95±1.4 84.06±1.6 86.21±1.0 85.37±0.9
Blankertz 71.10±0.8 70.64±0.6 72.19±0.9 76.23±1.4 70.59±1.7 69.50±1.8 74.41±1.1 73.66±0.9 72.72±0.9 75.40±1.3 73.21±0.8 73.43±1.1 76.33±0.7
2014002 72.86±0.4 70.57±1.4 74.34±0.8 71.31±0.8 72.67±0.4 74.14±0.8 72.84±1.4 74.09±0.6 73.99±1.1 74.61±1.1 73.01±0.6 73.90±0.7 77.17±0.8
OpenBMI 73.33±0.1 68.98±0.4 73.27±0.2 60.34±0.6 73.18±0.4 70.05±0.1 70.05±0.1 73.25±0.8 72.79±0.7 72.77±0.3 68.89±0.7 70.03±0.5 73.61±0.2
Average 73.65 71.12 73.23 71.60 72.58 72.71 72.84 74.51 73.94 75.37 71.99 74.30 76.67
Further, we conducted paired $t$-tests for each comparison and applied Benjamini-Hochberg False Discovery Rate correction to control for multiple comparisons [64]. The adjusted $p$-values are summarized in Table V, where values below 0.05 are highlighted in bold to indicate statistical significance. It is evident that the performance improvements of DBConformer over other competing models were almost always statistically significant.
TABLE V: Adjust $p$-values between DBConformer and twelve baseline models under CO, CV, and LOSO settings. Statistically significant results ($p<0.05$) are marked in bold.
Task DBConformer vs. BNCI2014001 BNCI2014004 Zhou2016 Blankerz2007 BNCI2014002 OpenBMI
CO EEGNet << 0.001 << 0.01 0.212 << 0.05 << 0.001 << 0.001
SCNN << 0.01 << 0.01 << 0.05 << 0.001 0.183 0.104
DCNN << 0.001 << 0.001 0.198 << 0.001 << 0.001 << 0.01
FBCNet << 0.001 << 0.01 << 0.001 << 0.001 << 0.001 << 0.001
ADFCNet << 0.001 << 0.01 << 0.01 << 0.01 << 0.01 0.081
CTNet << 0.01 0.052 0.054 << 0.001 << 0.01 << 0.01
EEG Conformer 0.052 << 0.01 << 0.01 << 0.01 << 0.05 << 0.051
MSCFormer << 0.001 0.051 << 0.05 0.13 << 0.001 << 0.05
TMSA-Net << 0.01 << 0.01 << 0.01 << 0.01 << 0.01 0.09
MSVTNet << 0.05 << 0.01 << 0.01 0.073 0.136 << 0.01
SlimSeiz << 0.001 << 0.05 << 0.05 << 0.001 0.116 << 0.001
IFNet << 0.05 << 0.05 0.158 << 0.05 0.128 << 0.001
CV EEGNet << 0.001 << 0.001 << 0.01 << 0.001 << 0.001 << 0.001
SCNN << 0.001 << 0.001 << 0.001 << 0.001 0.16 0.053
DCNN << 0.001 << 0.001 << 0.001 << 0.001 << 0.001 << 0.001
FBCNet << 0.001 << 0.001 << 0.001 << 0.001 << 0.001 << 0.001
ADFCNet << 0.001 << 0.01 << 0.001 << 0.001 << 0.001 0.132
CTNet << 0.001 << 0.01 << 0.001 << 0.01 << 0.001 << 0.001
EEG Conformer 0.08 << 0.01 << 0.001 << 0.01 0.099 << 0.05
MSCFormer << 0.01 << 0.05 << 0.01 0.089 0.063 << 0.05
TMSA-Net << 0.01 << 0.01 << 0.01 << 0.001 0.075 0.055
MSVTNet << 0.05 0.158 << 0.01 0.094 0.051 << 0.01
SlimSeiz << 0.001 << 0.01 << 0.01 << 0.001 << 0.01 << 0.001
IFNet << 0.05 << 0.01 << 0.001 << 0.05 0.064 << 0.01
LOSO EEGNet << 0.001 << 0.05 0.082 << 0.001 << 0.001 << 0.05
SCNN << 0.001 << 0.001 << 0.01 << 0.001 << 0.001 << 0.001
DCNN << 0.001 << 0.001 0.094 << 0.001 << 0.01 << 0.05
FBCNet << 0.001 << 0.001 << 0.001 0.172 << 0.001 << 0.001
ADFCNet << 0.001 << 0.001 << 0.01 << 0.001 << 0.001 0.055
CTNet << 0.001 << 0.001 0.082 << 0.001 << 0.01 << 0.001
EEG Conformer << 0.001 << 0.001 << 0.05 << 0.05 << 0.01 << 0.001
MSCFormer << 0.001 << 0.001 0.172 << 0.01 << 0.01 << 0.05
TMSA-Net 0.059 << 0.001 << 0.05 << 0.001 << 0.05 0.161
MSVTNet 0.14 << 0.001 0.182 0.076 0.054 << 0.05
SlimSeiz << 0.001 << 0.001 << 0.05 << 0.01 << 0.001 << 0.001
IFNet << 0.001 << 0.001 << 0.05 << 0.001 << 0.05 << 0.001
@@ IV-E Results on Seizure Detection
For seizure detection, we compared the proposed DBConformer with seven baseline models: EEGNet, SCNN, DCNN, ADFCNN, EEG Conformer, SlimSeiz, and EEGWaveNet. The first five models were originally designed for tasks such as MI and emotion recognition, rather than seizure detection. But they are general for all EEG decoding tasks; thus, we re-implemented and adapted them to seizure detection. SlimSeiz [30] is a CNN-Mamba hybrid model for seizure prediction that can be easily implemented in the seizure detection task. EEGWaveNet [26] is a multi-scale depthwise CNN specifically developed for seizure detection, serving as a competitive task-specific benchmark. Table VI summarizes the average AUCs, F1 scores, and BCAs of DBConformer and seven baseline models on the CHSZ and NICU datasets. Observe that:
@@
IV-F Results on SSVEP Classification
To further evaluate the generalizability of DBConformer, we conducted experiments on the Nakanishi2015 SSVEP dataset. As shown in Table VII, DBConformer consistently outperformed recent CNN and CNN-Transformer baselines, including Conformer, MSCFormer, TMSA-Net, and IFNet. In particular, DBConformer achieved the highest average accuracy of 87.59%, surpassing the second-best IFNet by a substantial margin, highlighting its adaptability across diverse EEG paradigms.
TABLE VII: Average classification accuracies (%) on Nakanishi2015 dataset under the CO setting. The best average performance is marked in bold, and the second-best by an underline.
Subject Conformer MSCFormer TMSA-Net IFNet DBConformer
S1 12.03 13.89 46.30 56.48 62.04
S2 8.33 12.96 24.07 32.41 46.30
S3 8.33 13.89 57.41 70.37 88.89
S4 30.55 14.81 78.70 93.52 96.29
S5 20.37 22.22 92.59 98.15 100.00
S6 35.18 15.74 87.04 92.59 100.00
S7 15.74 8.33 85.19 93.52 100.00
S8 16.66 19.44 91.67 100.00 100.00
S9 32.41 20.37 88.89 91.67 96.29
S10 10.18 11.11 54.63 65.74 86.11
Average 18.98 15.28 70.65 79.44 87.59
IV-G Ablation Study
To further validate the effectiveness of DBConformer’s key components, we conducted ablation studies on the parallel spatio-temporal design, positional encoding, and the channel attention module. Experiments were performed on six MI datasets under three settings, shown in Fig. 4. Observe that:
  • Parallel Spatio-Temporal Modeling. We compared the full DBConformer with its T-Conformer-only variant. Adding the S-Conformer branch consistently improved performance across all settings. Results confirmed the effectiveness of incorporating spatial features via the auxiliary S-Conformer, which complemented the dominant temporal modeling of T-Conformer.
  • Positional Encoding. Disabling the learnable positional embeddings for both temporal and spatial tokens resulted in performance drops under all settings. This validated the importance of explicit temporal/channel ordering when modeling patch and channel tokens.
  • Channel Attention Module. Replacing the channel attention module with naïve mean pooling degraded performance, confirming that adaptively reweighting EEG channels benefited enhancing spatial features.